Section 18

Teaching and Popularization of Mathematics

In case of several authors, Invited Speakers are marked with a ^{∗}.

George E. Andrews: Mathematics Education: Reform or Renewal? III 719

Michèle Artigue: De la Comprehension des Processus d’Apprentissage a la Conception de Processus d’Enseignement . . . . . . III 723

Maria G. Bartolini Bussi: Drawing Instruments: Theories and Practices from History to Didactics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 735

Miguel de Guzmán^{∗}, Bernard R. Hodgson^{∗}, Aline Robert^{∗} and Vinicio Villani^{∗}: Difficulties in the Passage from Secondary to Tertiary Education . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 747

D. J. Lewis: Mathematics Instruction in the Twenty-first Century . . III 763

Mogens Niss: Aspects of the Nature and State of Research in Mathematics Education . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 767

David A. Smith: Renewal in Collegiate Mathematics Education . . . . . III 777

Documenta Mathematica · Extra Volume ICM 1998 · III ·

718

Documenta Mathematica · Extra Volume ICM 1998 · III ·